APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday April 12, 2005 (MJD 53473) ---=== OBSERVING TEAM ===--- Scot Kleinman Atsuko Nitta Mike Harvanek (swing) Guest Appearances: Dan Long Jurek Krzesinski Howard Brewington John Barentine Steph Snedden (phone) Some good food! ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- 7 plates done - no major problems. One instance of low NLAMPS in a B1 arc. We tried to setup the DIMM, but the finder and CCD were so poorly mis-aligned, that we could not get Polaris on the CCD without spending lots of time trying. We left the telescope in the morning with French. ---=== OBSERVING LOG ===--- Afternoon: ---------- Spectro collimation was adjusted due to warmer temperatures forecast for tonight. Details are in the Problems section. No problems to report. Night: ------ non-photometric at the start, so we start with spectroscopy - cart 1 plate 2008. No guidestars, but it was also bright out... fk5 worked fine. 04:48Z - cartridge done in 4 exposures of 4900s total with a mid-calib. on to cart 3 plate 1695. GSOTGTF. 06:17Z - done in 3 exposures (3700s) on to cart 2 plate 2108. GSOTGTF. 07:06Z skies are clearing up, but weather map still looks iffy... 07:22Z: finished in 3 exposures(2300s) - on to cart 6 plate 1627. FK5 needed. The preCalib arc had bad B1 NLAMPS - we took a midCalib. 08:40Z: finished in 3 exposures (2800s) - on to cart 7 plate 1978. GSOTGTF. 09:54Z: finished in 3 exposures (2400s) with a requested postCalib. On to cart 4 plate 1579. GSOTGTF. 10:58Z finished in 3 exposures (2300s) - on to cart 9 plate 1692. GSOTGTF. To try to finish in time, we went straight to science exposures and did calibs as a set of postCalibs. We could not apoplot any fibers from this plate, getting errors like: IDL 5.5a> apoplot, 1692, 599, nsmooth=5 ; r=17.9168 APOPLOT: Selecting MJD=53473 (override this with MJD keyword) MRDFITS: File access error % Expression must be a structure in this context: TSET. % Execution halted at: TRACESET2XY 47 /local/ups/prd/Linux/idlutils/v4_10_7/pro/trace/traceset2xy.pro % APOPLOT1 182 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % APOPLOT 447 Is this due to having no valid preCalibs (we had dayshift test calibs) or is something else wrong? Finished (already existing data) in 2 additional exposures of 1520s. 12:00Z: French wanted the telescope to run some tests, so we left it with him. endNight finishing as this log is sent out. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 5246 23:46Z 00:03Z 100 O -92.92 -88.61 43 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y (but could not plot fibers from plate 1692) UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 31976, plate -9999 ------- 23:40 31976 0.0 bias ----- sequence 31977, plate 1692 ------- 00:07 31977 10.0 flat collimation 00:10 31978 2.0 arc collimation 00:14 31979 2.1 arc collimation 00:17 31980 2.1 arc collimation 00:31 31981 10.0 flat collimation check 00:34 31982 2.0 arc collimation check 00:37 31983 2.1 arc collimation check 00:39 31984 2.1 arc collimation check ----- sequence 31985, plate -9999 ------- 01:36 31985 0.0 bias 02:09 31986 0.0 bias ----- sequence 31987, plate 2008 ------- 02:47 31987 10.0 flat 02:50 31988 2.0 arc 03:15 31989 1200.1 target 03:46 31990 1200.1 target 04:14 31991 1500.1 target 04:18 31992 10.0 flat 04:21 31993 2.0 arc 04:42 31994 1000.1 target DONE ----- sequence 31995, plate 1695 ------- 04:58 31995 10.0 flat 05:01 31996 2.0 arc 05:26 31997 1200.1 target 05:54 31998 1500.1 target 06:14 31999 1000.1 target DONE ----- sequence 32000, plate 2108 ------- 06:25 32000 10.0 flat 06:28 32001 2.0 arc 06:47 32002 900.1 target 07:02 32003 700.1 target 07:19 32004 700.1 target DONE ----- sequence 32005, plate 1627 ------- 07:32 32005 10.0 flat 07:35 32006 2.0 arc 07:51 32007 700.1 target 07:54 32008 2.0 arc 08:13 32009 900.1 target 08:36 32010 1200.1 target DONE ----- sequence 32011, plate 1978 ------- 08:48 32011 10.0 flat 08:53 32012 2.0 arc 09:18 32013 1200.1 target 09:31 32014 600.1 target 09:44 32015 600.2 target 09:47 32016 10.0 flat 09:50 32017 2.0 arc DONE ----- sequence 32018, plate 1579 ------- 10:02 32018 10.0 flat 10:05 32019 2.0 arc 10:24 32020 900.1 target 10:39 32021 700.1 target 10:54 32022 700.1 target DONE ----- sequence 32023, plate 1692 ------- 11:19 32023 720.1 target 11:35 32024 800.1 target 11:38 32025 10.0 flat 11:41 32026 2.0 arc DONE ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:53Z 1 2008 98.00 0.0017 58.95 0.0053 257.77 -0.0082 1.000250 05:42Z 3 0 32.72 0.0008 66.45 0.0046 207.62 0.0132 1.000170 06:32Z 2 2108 236.71 0.0068 78.78 0.0043 63.60 0.0355 1.000170 07:39Z 6 1627 -45.05 -0.0007 56.96 0.0042 143.01 -0.0188 1.000170 09:10Z 7 1978 271.33 0.0027 63.59 0.0035 106.95 -0.0145 1.000080 10:07Z 4 1579 285.54 0.0036 81.92 0.0036 110.13 0.0318 0.999980 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL8458 Stays: JL8459 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 02:53Z 1 2008 1.00025 -2468 -888 0 98 58.9 7.6 10 32 gdr 2.0 05:42Z 3 1695 1.00017 -1678 -277 -25 33 66.4 7.4 1 132 gdr 1.2 06:32Z 2 2108 1.00017 -1678 -270 0 237 78.8 7.2 4 120 gdr 1.0 07:39Z 6 1627 1.00017 -1678 -189 40 -45 57.0 7.5 2 142 gdr 1.2 09:10Z 7 1978 1.00008 -790 465 -25 271 63.6 6.8 2 190 gdr 1.2 10:08Z 4 1579 0.99998 198 1255 -25 285 81.9 7.2 0 236 gdr 1.2 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:24Z 49 4 2 338 (NNW) 1673 - 23:57Z 49 2 7 297 (WNW) 1688 - 00:29Z 48 1 8 328 (NNW) 1701 - 01:01Z 46 1 6 1 (N) 1934 - 01:33Z 46 2 4 357 (NNW) 1844 - partly cloudy 02:06Z 45 1 5 12 (NNE) 2003 - 02:38Z 45 2 7 4 (N) 1954 - 03:10Z 44 2 13 42 (NE) 1991 - 03:41Z 45 4 12 63 (ENE) 1888 - 04:12Z 45 4 12 72 (ENE) 1962 - 04:43Z 44 4 10 81 (E) 1921 - 05:15Z 45 4 4 88 (E) 2118 - 05:47Z 44 1 2 115 (ESE) 2119 - 06:20Z 44 3 5 114 (ESE) 1839 - 06:50Z 43 2 11 137 (SE) 1848 - 07:20Z 43 2 8 127 (SE) 1835 - 07:53Z 46 4 1 178 (S) 1294 - 08:25Z 46 1 0 146 (SE) 1624 - 08:56Z 44 1 5 258 (WSW) 1351 - 09:27Z 44 2 1 184 (S) 1784 - 09:57Z 44 3 1 111 (ESE) 1600 - 10:27Z 44 2 1 191 (S) 1684 - 10:58Z 44 2 2 109 (ESE) 1562 - 11:29Z 44 1 6 93 (E) 1810 - 11:59Z 45 -1 8 87 (E) 3020 - ---=== TELESCOPE STATUS ===--- 01:30Z Enclosure doors opened, fans and blowers on, louvers opened. 01:55Z Enclosure off. 12:00Z: French wanted the telescope for some tests, so we left it with him. ---=== SOFTWARE USED ===--- IOP/SOP: v3_143_1 Watcher: v2_30_0 MCP: v5_26_0 TPM: tpm_v2_47_0 AstroDa: v14_47 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v1_98 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5760 0.5770 Axial B 0.5890 0.5880 Axial C 0.9450 0.9460 Trans D -8.2040 -8.1960 Lateral E 10.0457 10.0330 Lateral F 11.6078 11.5951 GALILS Commanded: 4300. -1900. 2850. -13300. -7600. -7650. Actual: 4299. -1917. 2868. -13303. -7614. -7660. SETMIR VALUES PriDesOrient: 0.00 4.57 1.18 64.25 -249.76 PriOrient: 0.00 4.56 0.88 65.00 -249.58 SECONDARY: ---------- Focus: 0.00 Air Temp.: 12.1 Alt.: 30.188289 MIGS TONIGHT NOMINAL Axial A 1.5700 1.5680 Axial B 1.0470 1.0460 Axial C 1.1620 1.1590 Trans D -0.2100 -0.2020 GALILS Commanded: 1619113. 1604928. 1528418. -13800. -9950. Actual: 1619188. 1604973. 1528571. -13820. -9964. SETMIR VALUES SecDesOrient: 1257.00 27.97 -10.00 -90.00 257.89 SecOrient: 1256.93 28.00 -10.00 -89.94 257.48 ---=== PROBLEMS IN DETAIL ===--- Spectro collimation details: ---------------------------- Initial sp1 Coll_motor_A -4701 Coll_motor_B -4707 Coll_motor_C -3322 Initial sp2 Coll_motor_A -1701 Coll_motor_B -3107 Coll_motor_C -4401 im> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 7.30357 tpm_TM_TelairAvg.VALA 9.3875 Adjustment performed due to warmer temperatures forecast for tonight. Spectro collimation set for +6 C. sp1> mechPiston 8000 sp2> mechPiston 8000 Final sp1 Coll_motor_A 3300 Coll_motor_B 3295 Coll_motor_C 4678 Final sp2 Coll_motor_A 6300 Coll_motor_B 4898 Coll_motor_C 3600 Final collimate output COLLIMATE: Mean offset = 0.18 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Mean offset = 0.07 pix COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Mean offset = -0.24 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2152 steps for spectro-1 COLLIMATE: Mean offset = -0.15 pix COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 1360 steps for spectro-2 Watcher errors: --------------- We got lots of: rotator (addoffset: Max accl. for rotator exceeded: tonight.